OpenStudy (anonymous):

Each year on the same day, Susan deposits $175 into a savings account that earns simple interest at a rate of 3.5%. She makes no withdrawals. How much interest does Susan’s account earn after 5 years? A $18.38 B $30.65 C $61.25 D $91.89

OpenStudy (qwertty123):

Susan is putting money into a checking account. Let y represent the total amount of money in the account (in dollars). Let x represent the number of weeks Susan has been adding money. Suppose that x and y are related by the equation = + 30 x 350 y

OpenStudy (anonymous):

not really.. :/

OpenStudy (mathstudent55):

Do you know the formula for simple interest?

OpenStudy (study312):

correct me if I am wrong but I think its 175 • 5 = 875 875 • 0.035 = 30.625 which would be B

OpenStudy (study312):

@mathstudent55 is what I just posted right?

OpenStudy (mathstudent55):

I don't think that is it. You treated the problem as if a deposit of $875 were made originally, and then the question is how much interest accrued at the end of 5 years.

OpenStudy (mathstudent55):

I don't think that is it. You treated the problem as if a deposit of $875 were made originally, and then you calculated how much interest was earned for only 1 year. Simple interest means that there is no interest earned on interest. (For earning interest on interest you need compound interest, not in this problem.) Once again, simple interest only earns interest on the principal. In this problem, the principal changes every year, as a new $175 deposit is made. The first $175 deposited will be there for 5 years. Calculate 3.5% of $175 for 5 years. The next $175 deposit at the end of the first year will be there for another 4 years. Calculate 3.5% of $175 for 4 years. The next $175 deposit at the end of the second year will be there for another 3 years. Calculate 3.5% of $175 for 3 years. The next $175 deposit at the end of the third year will be there for another 2 years. Calculate 3.5% of $175 for 2 years. The next $175 deposit at the end of the fourth year will be there for another 1 year. Calculate 3.5% of $175 for 1 years. Then add all these amounts.

OpenStudy (study312):

I'll give you a medal just for taking 10 min to write that -clap clap clap clap-

OpenStudy (study312):

bravo...bravo

OpenStudy (qwertty123):

YEAH!!!!!

OpenStudy (anonymous):

I did what you said @mathstudent55 and I got 93.7358515625

OpenStudy (mathstudent55):

I = Prt interest = principal * rate * time In our problem, the rate is 3.5%, which is 0.035 as a decimal The principal is $175 for each deposit. The time is 5 years for the first deposit, 4 years for the second deposit, etc. The first deposit of $175 is earning interest for 5 years at 3.5% per year. \(I_5 = $175 \times 0.035 \times 5 = $30.63\) The second deposit of $175 is earning interest for 4 years at 3.5% per year. \(I_4 = $175 \times 0.035 \times 4 = $24.50\) The third deposit of $175 is earning interest for 3 years at 3.5% per year. \(I_3 = $175 \times 0.035 \times 3 = 18.38\) The fourth deposit of $175 is earning interest for 2 years at 3.5% per year. Total interest: $91.89 \(I_2 = $175 \times 0.035 \times 2 = $12.25\) The fifth deposit of $175 is earning interest for 1 year at 3.5% per year. \(I_1 = $175 \times 0.035 \times = $6.13\) Add all the amounts of interest earned to find the total interest earned.

OpenStudy (mathstudent55):

I got a total of $91.89

OpenStudy (anonymous):

Oh I see what you did

OpenStudy (mathstudent55):

I = Prt interest = principal * rate * time In our problem, the rate is 3.5%, which is 0.035 as a decimal The principal is $175 for each deposit. The time is 5 years for the first deposit, 4 years for the second deposit, etc. The first deposit of $175 is earning interest for 5 years at 3.5% per year. \(I_5=$175×0.035×5=$30.63\) The second deposit of $175 is earning interest for 4 years at 3.5% per year. \(I_4=$175×0.035×4=$24.50\) The third deposit of $175 is earning interest for 3 years at 3.5% per year. \(I_3=$175×0.035×3=18.38\) The fourth deposit of $175 is earning interest for 2 years at 3.5% per year. \(I_4=$175×0.035×2= $12.25\) The fifth deposit of $175 is earning interest for 1 year at 3.5% per year. \(I_1=$175×0.035×=$6.13\) Total interest: $91.89

OpenStudy (mathstudent55):

I just posted my answer again because the previous one came out in the incorrect order.